OpenStudy (anonymous):

I need help simplifying the square root of -48. The choices are -4 sqrt of 3, 4 sqrt of -3, 4i sqrt of 3, 4 sqrt of 3i please help

OpenStudy (anonymous):

@Ankh

OpenStudy (anonymous):

Thanks again!!

OpenStudy (anonymous):

You want the steps for these?

OpenStudy (anonymous):

I think it is 4i sq root of 3 cause you split -48 as -1x 16x 3 and take root of -1 it is I, root 16 is 4 left over is sq root of 3

OpenStudy (anonymous):

@mila1628 \[\sqrt{-1} \times \sqrt{48}\] \[i \times \sqrt{16\times3}\] \[i4\sqrt{3}\] Yes mila, you are correct.